40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 5457     Entry time: Mon Sep 19 12:23:30 2011
Author: Paul 
Type: Update 
Category: SUS 
Subject: ITMY and SRM oplev beam size reduced + next steps 

I replaced the lenses that were there with a -150mm lens followed by a +250mm lens. This gave a significantly reduced beam size at the QPDs. With the beam analyzer up and running it should be possible to optimize this later this afternoon. Next I will remove the SRM QPD from the path and make measurements of the beam spot position movement and corresponding OSEM values for different DC mirror offsets. I will then repeat the process for ITMY.

ELOG V3.1.3-